# MRIR: Integrating Multimodal Insights for Diffusion-based Realistic Image Restoration

### v1

@@ Abstract
Realistic image restoration is a crucial task in computer vision, and the use of diffusion-based models for image restoration has garnered significant attention due to their ability to produce realistic results. However, the quality of the generated images is still a significant challenge due to the severity of image degradation and the uncontrollability of the diffusion model. In this work, we delve into the potential of utilizing pre-trained stable diffusion for image restoration and propose MRIR, a diffusion-based restoration method with multimodal insights. Specifically, we explore the problem from two perspectives: textual level and visual level. For the textual level, we harness the power of the pre-trained multimodal large language model to infer meaningful semantic information from low-quality images. Furthermore, we employ the CLIP image encoder with a designed Refine Layer to capture image details as a supplement. For the visual level, we mainly focus on the pixel level control. Thus, we utilize a Pixel-level Processor and ControlNet to control spatial structures. Finally, we integrate the aforementioned control information into the denoising U-Net using multi-level attention mechanisms and realize controllable image restoration with multimodal insights. The qualitative and quantitative results demonstrate our method’s superiority over other state-of-the-art methods on both synthetic and real-world datasets.
@@ I Introduction
Image restoration (IR) has always been a classic problem in digital image processing, aiming to convert low-quality (LQ) images into high-quality (HQ) and clear images. Typical image restoration tasks include super-resolution (SR) [1, 2, 3], deblurring [4, 5, 6], denoising [7, 8], color enhancement [9, 10], inpainting [11, 12, 13], compression artifact removal [14, 15], etc. Early research often focused on studying image restoration problems under the assumption of a known linear degradation type present in the image. However, real-world low-quality images often contain complex and diverse degradation types, making these methods less effective in practical applications. Therefore, BSRGAN [16] and Real-ESRGAN [17] proposed new degradation synthesis methods to model the degradation process and achieve realistic image restoration (real-IR). As image restoration techniques have advanced, there is a significant increase in the expectation of the perceptual quality of restored images. The over-smoothed results produced by the methods mentioned above fail to meet these expectations. To address this, generative adversarial networks (GANs) have been employed to solve the real-IR problem [18, 19, 20, 21]. By using adversarial loss, supervised image restoration models generate perceptually realistic details. However, GAN-based real-IR methods often introduce unnatural visual artifacts [20, 21].
@@
Recently, denoising diffusion probabilistic models (DDPMs) [22] have achieved significant success in the field of image generation, surpassing traditional generative adversarial networks (GANs) in various downstream tasks [23]. As a result, researchers have started exploring the use of DDPMs for image restoration, incorporating high-quality generative priors from diffusion models into image restoration tasks to generate more realistic images[24, 25, 26, 27]. Alongside the success of large-scale pre-trained text-to-image (T2I) models, like Stable Diffusion (SD) trained on datasets containing over 5 billion image-text pairs, which provide rich natural image priors, serving as a vast library of textures and structures, some methods [28, 29, 30, 31, 32] have attempted blind image super-resolution using pre-trained T2I diffusion models. However, these methods have not fully exploited the potential of T2I models, particularly in obtaining high-quality textual prompts from low-quality images. Furthermore, with the development of multi-modal techniques, large language models (LLMs) or multimodal large language models (MLLMs) have played a significant role in multi-modal high-quality image generation and editing tasks [33, 34, 35, 36]. However, applying LLMs or MLLMs to image restoration tasks is still limited [37, 34].
In this paper, we delve into the exploration of harnessing the generative potential of pre-trained T2I models for achieving realistic image restoration. We approach the problem from both textual level and visual level perspectives, proposing a multimodal-perception realistic image restoration model (MRIR). For the textual level, to obtain high-quality textual representations from LQ images, we introduce a multimodal large language model to predict the semantic prompt of restored images from LQ inputs. Additionally, to ensure fidelity to the original image during the restoration process, we design an image embedding branch that collaborates with the text predicted by MLLM, achieving a balance between fidelity and realism. Regarding the visual level, we focus on pixel control and employ a Pixel-level Processor to extract clear features from the degraded image to control the restoration process. Finally, we employ attention mechanisms to integrate pixel control, text embedding, and image embedding to the denoising U-Net of the diffusion model. All modules collaborate to generate structurally accurate and detail-rich restored images.
@@
  • We propose a diffusion-based realistic image restoration model that incorporates dual-perception of textual and visual level information.
  • We utilize MLLM and refine layers to obtain effective textual control. We employ a Pixel-level Processor and multi-layer supervision to get clear pixel control.
  • We integrate the control information with the diffusion model using multi-layer attention mechanisms. Our model demonstrates advanced performance on multiple datasets, as measured by various image quality metrics.
@@ II-A Realistic Image Restoration
The goal of realistic image restoration is to transform the LQ images with degradation artifacts to HQ and photo-realistic ones. Early research focused on image restoration of single degradation types, such as super-resolution [1, 2, 3, 38, 39, 40], denoising [7, 8, 41], deblurring [4, 5, 6], and color enhancement [9, 10]. Some researchers are dedicated to studying baseline models for image restoration to handle various types of degradation[42, 43, 44], but training on specific datasets is required for each degradation type. Overall, these methods often rely on specific degradation types, while real-world degradation types are typically complex and diverse, making it challenging to apply these methods in practical scenarios. Therefore, some methods [16, 17, 45] have made progress in the field of realistic image restoration by modeling the degradation process of images to simulate real-world image degradation. With the development of Generative Adversarial Networks, some studies[16, 17, 18, 19, 20, 21] have utilized pre-trained GANs to improve the super-resolution process, reducing the smoothness of generated images and enhancing the richness of texture details. Specifically, BSRGAN [16] utilizes a randomly selected degradation strategy to enhance the richness of degradation, while Real-ESRGAN [17] employs a high-order degradation modeling process. DASR [18] proposes an adaptive super-resolution network that estimates the degradation for each input image. FeMaSR [19] utilizes a pre-trained VQGAN network for feature matching between low-resolution features and the distortion-free high-resolution prior, resulting in more realistic and less artifact-prone image generation. However, GAN-based real-IR methods often introduce unnatural visual artifacts [20, 21]. Therefore, in recent research, there is increasing interest in utilizing more advanced pre-trained generative models, such as denoising diffusion models [22, 23, 46], for image restoration to generate more realistic images.
@@ II-B Diffusion-Based Image Restoration
With the development of diffusion models, several attempts [24, 25, 26] have been made to utilize DDPM [22] for addressing image super-resolution problems. However, these methods often rely on simple downsampling linear degradation, making it challenging to apply them effectively in practical restoration scenarios. The assumption of known linear image degradation limits their practical application in complex scenes. Concurrently, other studies [28, 29, 30, 31, 32], have employed powerful pre-trained T2I models, such as stable diffusion [47], to tackle real-world image restoration problems. Trained on billions of image-text pairs, these models possess rich image priors that are crucial for real-world image restoration. StableSR [28] finetunes the SD model by training a time-aware encoder and an additional CFW module [48] to balance fidelity and perceptual quality, enabling blind image restoration. DiffBIR [29] employs a two-stage strategy to address real-IR problems. It first reconstructs an image as an initial estimate and then employs SD priors to enhance image details. The aforementioned methods rely solely on images as conditions to activate the generation capability of T2I models. In contrast, PASD [30] takes it a step further by utilizing high-level models, but the acquisition of high-level information relies on low-quality images, leading to a substantial amount of inaccurate information. SeeSR [31] trains a dedicated RAM model [49] for assistance in restoration, while CoSeR [32] trains a dedicated cognitive encoder for object categorization. However, both methods require separate training, and their restoration performance is dependent on the additional trained RAM and cognitive encoder. What’s more, they perform better when dealing with images containing explicitly categorized objects, but struggle with images lacking clear category information. Therefore, further exploration is needed on how to introduce high-quality information from low-quality images to more effectively leverage the potential of pre-trained T2I models for assisting real-IR.
@@ II-C MLLM Asssited Image Restoration
With the rapid development of text-to-image techniques, multimodal image generation has attracted increasing attention. LLMs and MLLMs have provided a range of solutions for high-quality multimodal image generation and editing tasks [33, 34, 35, 36]. However, only a few works [37, 34, 50] have extended LLMs or MLLMs to image restoration tasks. Clarity ChatGPT [37] combines ChatGPT with various pre-trained image restoration models, utilizing the rich prior reasoning capabilities of LLMs/MLLMs to determine, select, and invoke appropriate image restoration base models. LLMRA [34] utilizes pre-trained MLLMs to generate descriptive textual descriptions of the input degraded images, encompassing detailed textual descriptions of low-level features in the degraded images. These descriptions are encoded as text features and serve as guiding conditions that are input into a transformer network responsible for image restoration. [50] employs MLLMs to simultaneously take input images and chat-style user requests, generating language dialogue responses and an implicit guiding prompt, which are input into the cross-attention module of the generation network through a mapping network to guide image inpainting. However, these methods are only simple applications of MLLM and are not combined with the diffusion model to handle more realistic image restoration. SUPIR [51] applies text guidance from MLLMs in diffusion models. It’s the first attempt to combine the MLLM and diffusion to realize image restoration, but the method applies the MLLM to describe the images based on the pre-restored images and does not fully exploit the inferential capabilities of MLLMs for predicting the content of the restored images.
@@ III Method
III-A Motivation and Overview
Motivation. Motivated by the remarkable success of ControlNet [52] in controllable image generation, researchers have turned their attention to employing the ControlNet architecture with the pre-trained T2I models for realistic image restoration. Thus, the quality of the generated images by the diffusion model is heavily influenced by two crucial factors: prompts and control conditions of ControlNet. Consequently, extensive research has been conducted from these two perspectives.
In terms of the textual control conditions, various approaches have been investigated. We conduct a preliminary experiment with different textual prompts. We use the PASD [30] as a baseline and employ null prompts, LQ prompts (captions generated from LQ images), and GT prompts (captions generated from GT images) to guide image restoration individually. Fig 1 illustrates the impact of different textual information on the restoration results. As shown in Fig 1, compared with generation with null prompts, incorporating textual prompts proves beneficial for generating detailed image content, and the richness of generated details increases with more detailed textual descriptions. However, due to the degradation in the LQ images, directly utilizing the captions derived from these images may lead to semantic errors. As depicted in Fig 1 (b), the prompt extracted from the LQ image results in an incorrect description, leading the T2I model to erroneously generate a plane. In fact, obtaining GT prompts for realistic image restoration is not feasible. Consequently, we aim to infer accurate and detailed textual descriptions from LQ images to guide the restoration process. Leveraging the remarkable success of MLLMs, we employ the MLLM to predict the semantic content. Taking the LQ image in Fig 1 as an example, the prompt predicted by the MLLM is as follows: “A field of ripe wheat stands against a clear blue sky, creating a picturesque sight. The golden wheat glows under the sunlight, while the serene blue sky provides a beautiful backdrop.” The predicted prompt possesses accurate semantics and rich descriptive details. When used to guide the PASD restoration, it achieves comparable or even superior restoration result to the result obtained using the GT prompt, as shown in Fig 1 (d).
In terms of pixel-level control, we aim to extract clear features from LQ images to exert control over the generated image. Therefore, we employ preprocessing techniques to obtain clear features and utilize ControlNet to regulate the diffusion process.
In summary, we construct our image restoration model by considering the two aforementioned perspectives. Firstly, in the context of textual control conditions, our objective is to infer semantically accurate and detailed textual descriptions from low-quality images, facilitating precise control over the image generation process. Secondly, for the pixel control, we aim to exploit the latent clear features from the low-quality images to effectively guide the restoration process.

### v2

# SSP-IR: Semantic and Structure Priors for Diffusion-based Realistic Image Restoration

@@ Abstract
Realistic image restoration is a crucial task in computer vision, and diffusion-based models for image restoration have garnered significant attention due to their ability to produce realistic results. Restoration can be seen as a controllable generation conditioning on priors. However, due to the severity of image degradation, existing diffusion-based restoration methods cannot fully exploit priors from low-quality images and still have many challenges in perceptual quality, semantic fidelity, and structure accuracy. Based on the challenges, we introduce a novel image restoration method, SSP-IR. Our approach aims to fully exploit semantic and structure priors from low-quality images to guide the diffusion model in generating semantically faithful and structurally accurate natural restoration results. Specifically, we integrate the visual comprehension capabilities of Multimodal Large Language Models (explicit) and the visual representations of the original image (implicit) to acquire accurate semantic prior. To extract degradation-independent structure prior, we introduce a Processor with RGB and FFT constraints to extract structure prior from the low-quality images, guiding the diffusion model and preventing the generation of unreasonable artifacts. Lastly, we employ a multi-level attention mechanism to integrate the acquired semantic and structure priors. The qualitative and quantitative results demonstrate that our method outperforms other state-of-the-art methods overall on both synthetic and real-world datasets. Our project page is https://zyhrainbow.github.io/projects/SSP-IR.
@@ I Introduction
Image restoration (IR) has always been a classic problem in digital image processing, aiming to convert low-quality (LQ) images into high-quality (HQ) images. Typical image restoration tasks include super-resolution (SR) [1, 2, 3], deblurring [4, 5, 6], denoising [7, 8], color enhancement [9, 10], inpainting [11, 12, 13], and compression artifact removal[14, 15]. Early research often focused on studying image restoration problems under the assumption of a known linear degradation type present in the image. However, real-world low-quality images usually contain complex and diverse degradation types, making these methods less effective in practical applications. Therefore, BSRGAN[16] and Real-ESRGAN[17] have proposed new degradation synthesis methods to model the degradation process and achieve realistic image restoration (real-IR). As image restoration techniques have advanced, there is a significant increase in the expectation of the perceptual quality of restored images. The over-smoothed results produced by the abovementioned methods fail to meet these expectations. To address this problem, generative adversarial networks (GANs) have been employed to solve the real-IR problem [18, 19, 20, 21, 22]. By using adversarial loss, the restoration models can generate perceptually realistic results. However, GAN-based real-IR methods often introduce unnatural visual artifacts [20, 21].
Figure 1: Illustration of our method’s superiority compared with other diffusion-based methods. (a) shows our method’s good performance on the semantic, structure and perception quality. We calculate the CLIP similarity between the generated results and GT images to evaluate the semantic fidelity, SSIM to evaluate the structure accuracy and CLIPIQA to evaluate the perception quality. All metrics are calculated on RealSR dataset [23] and are the higher the better. Our method achieves a good balance about semantic-structure-perception. In the figure, $\checkmark$ indicates better performance and $\times$ indicates poorer performance. (b) shows the visual comparison compared with other diffusion-based methods.
@@
Recently, denoising diffusion probabilistic models (DDPMs) [24] have achieved remarkable success in the field of image generation, surpassing traditional generative adversarial networks (GANs) in various downstream tasks [25]. As a result, researchers have started exploring the application of DDPMs for image restoration, incorporating high-quality generative priors from diffusion models into image restoration tasks[26, 27, 28, 29, 30]. This approach aims to generate more realistic images without unnatural artifacts. Large-scale pre-trained text-to-image (T2I) models have gained great success. Particularly, stable diffusion (SD) which was trained on datasets containing over 5 billion image-text pairs, provides rich natural image priors and serves as a vast library of textures and structure. Some methods [31, 32, 33, 34, 35] have attempted to apply pre-trained stable diffusion to the blind image super-resolution task.
Unlike traditional image restoration approaches, diffusion-based methods adhere to the principle of “reconstruction as conditional generation”, aiming to introduce control conditions to guide the diffusion model. DiffBIR [32] trains a preprocessing model and introduces structure control to the diffusion model. However, this method requires two-stage training and lacks semantic prompts to assist the generation process. In addition, the preprocessing model model may lead to over-smoothing and reduce the structure accuracy. PASD [33] utilizes existing BLIP [36] to generate semantic prompts, yet the presence of artifacts in degraded images can lead to inaccurate extraction of semantic information as shown in Fig 3. SeeSR [34] trains a dedicated RAM model [37] to generate semantic tags, while CoSeR [35] trains a specialized cognitive encoder to generate object embedding. However, both SeeSR and CoSeR require separate training, and their performance depends on the additional training of RAM and cognitive encoders. They perform better when handling images with explicitly categorized objects but struggle with images lacking clear category information.
With the advancement of multimodal technologies, Large Language Models (LLMs) or Multimodal Large Language Models (MLLMs) have played crucial roles in multimodal high-quality image generation and editing tasks [38, 39, 40, 41]. SUPIR [42] pioneers the application of text guidance from MLLM in restoration models, attempting to integrate MLLM and diffusion for image restoration. However, this method relies on prompts from pre-restored images, requiring additional training to generate these pre-restored images. Furthermore, the prompts are based on pre-restored images, which may not be faithful to the input. Moreover, SUPIR demands substantial training resources. Thus, the challenge of effectively leveraging prior conditions in degraded images to generate semantically faithful, structurally accurate, and perceptually high-quality images is a critical consideration.
In this paper, we present an effective solution SSP-IR, which fully exploits the semantic and structure priors in the input image for conditional generation. Specifically, for Semantic Prior, we adopt an explicit-implicit collaboration strategy to obtain richer and more accurate semantic information. We first utilize a multimodal large language model to generate detailed textual descriptions, which are explicit and can be manually adjusted by users to tailor the restoration results to meet requirements. Additionally, to acquire semantic information that is more faithful to the original image, we extract the image embedding of the input image with CLIP image encoder. We also design a Refine Layer to mitigate the influence of degraded artifacts. The explicit and implicit semantic information interact with each other to achieve a balance between fidelity and realism. Regarding Structure Prior, we introduce a constrained Processor to obtain degradation-independent structure prior. Finally, we integrate semantic and structure priors through Prior-Guided Attention Module to the denoising U-Net, achieving high-quality restoration results. Fig 1 shows the superiority of our method in the quality of semantic, structure, and perception.
@@
  • We propose an effective one-stage realistic image restoration framework with semantic and structure priors to achieve high-quality restoration results.
  • We introduce an explicit-implicit collaboration strategy to obtain accurate and specific semantic information by integrating the visual understanding capabilities of MLLM and the intrinsic representation capabilities of images.
  • To ensure fidelity to input and avoid introducing degradation artifacts, we employ a preprocessing module with RGB and FFT constraints to extract degradation-independent structure features.
  • We employ a Prior-Guided Attention module to integrate the semantic and structure priors to the diffusion model. Qualitative and quantitative results on multiple datasets demonstrate the superiority of our method in terms of semantic, structure, and perception.
@@ II-A Realistic Image Restoration
The goal of realistic image restoration is to transform the LQ images with degradation artifacts to HQ and photo-realistic ones. Early research focused on image restoration of single degradation types, such as super-resolution [1, 2, 3, 43, 44, 45], denoising [7, 8, 46], deblurring [4, 5, 6], and color enhancement [9, 10]. Some researchers are dedicated to studying baseline models for image restoration to handle various types of degradation[47, 48, 49], but training on specific datasets is required for each degradation type. Overall, these methods often rely on specific degradation types, while real-world degradation types are typically complex and diverse, making it challenging to apply these methods in practical scenarios. Therefore, some methods [16, 17, 50] have made progress in the field of realistic image restoration by modeling the degradation process to simulate real-world image degradation.
With the development of Generative Adversarial Networks, some studies[16, 17, 18, 19, 20, 21] have utilized pre-trained GANs to improve the super-resolution process, reducing the smoothness of generated images and enhancing the richness of texture details. Specifically, BSRGAN [16] utilizes a randomly selected degradation strategy to enhance the richness of degradation, while Real-ESRGAN [17] employs a high-order degradation modeling process. DASR [18] proposes an adaptive super-resolution network that estimates the degradation for each input image. FeMaSR [19] utilizes a pre-trained VQGAN for feature matching between low-resolution features and the distortion-free high-resolution prior, resulting in more realistic and less artifact-prone image generation. However, GAN-based real-IR methods often introduce unnatural visual artifacts [20, 21]. Therefore, in recent research, there is increasing interest in utilizing more advanced pre-trained generative models, such as denoising diffusion models [24, 25, 51], for image restoration to generate more realistic images.
@@ II-B Diffusion-Based Image Restoration
With the development of diffusion models, several attempts [26, 27, 28] have been made to utilize DDPM [24] for addressing image super-resolution problems. However, these methods often rely on simple linear degradation, making it challenging to apply them effectively in practical restoration scenarios. The assumption of known linear image degradation limits their practical application in complex scenes.
Concurrently, other studies [31, 32, 33, 34, 35] have employed powerful pre-trained T2I models, such as stable diffusion [52], to tackle real-world image restoration problems. Trained on billions of image-text pairs, stable diffusion possesses rich image priors that are crucial for real-world image restoration. StableSR [31] finetunes the SD model by training a time-aware encoder and an additional CFW module [53] to balance fidelity and perceptual quality, enabling blind image restoration. DiffBIR [32] employs a two-stage strategy to address real-IR problems. It first reconstructs an image as an initial estimate and then employs SD priors to enhance image details. The aforementioned methods rely solely on images as conditions to activate the generation capability of T2I models. In contrast, PASD [33] takes it a step further by utilizing high-level models, but the acquisition of high-level information relies on low-quality images. SeeSR [34] trains a dedicated RAM model [37] for assistance in restoration, while CoSeR [35] trains a dedicated cognitive encoder for object categorization. However, both SeeSR and CoSeR require additional training, and their restoration performance is dependent on the additional trained RAM and cognitive encoder. Moreover, they perform better when dealing with images containing explicitly categorized objects, but struggle with images lacking clear category information.
In conclusion, existing diffusion-based methods still have certain limitations in the real-IR task. Therefore, further exploration is needed on how to introduce semantic and structure information from low-quality images to more effectively leverage the potential of pre-trained T2I models in assisting real-IR.
@@ II-C MLLM Asssited Image Restoration
With the rapid development of text-to-image techniques, multimodal image generation has attracted increasing attention. LLMs and MLLMs have provided a range of solutions for high-quality multimodal image generation and editing tasks [38, 39, 40, 41]. However, only a few works [54, 39, 55] have extended LLMs or MLLMs to image restoration tasks. Clarity ChatGPT [54] combines ChatGPT with various pre-trained image restoration models, utilizing the rich prior reasoning capabilities of LLMs/MLLMs to determine, select, and invoke appropriate image restoration base models. LLMRA [39] utilizes pre-trained MLLMs to generate detailed descriptions of the input degraded images. These descriptions are encoded as text features and serve as guiding conditions for transformer-based image restoration. [55] employs MLLMs to simultaneously process input images and chat-style user requests. It generates language dialogue responses as implicit guiding prompts, which are then used to guide image inpainting. However, these methods are only simple applications of MLLM and are not combined with the diffusion model to handle realistic image restoration.
SUPIR [42] applies text guidance from MLLMs in diffusion models. It combines the potential of MLLM and SD to realize image restoration, but the method relies on text descriptions generated on pre-restored images, which requires additional training to generate the pre-restored images. Moreover, the generated text descriptions are based on the pre-restored images, which may not be faithful to the inputs. In contrast, our approach directly utilizes the original image as input and leverages the visual understanding capability of MLLM to predict the content of the target restored image. This allows us to obtain text prompts that are more informative and faithful to the original image. In addition, we also utilize the image embedding of the input as a supplement, which is more conducive to generating the result faithful to the original image.
@@ III Method
Figure 2: The framework of the proposed method, which consists of semantic prior extraction module, structure prior extraction module and the denoising U-Net. Semantic prior extraction module consists of two branches and is used to generate explicit text embedding and implicit image embedding. Structure prior extraction module is used to generate degradation-independent structure control. The denoising U-Net integrates the text embedding, image embedding and strcuture control into the denoising process through the Prior-Guided Attention module.
In this section, we first describe the overall framework of our proposed SSP-IR. Then we describe how we get the semantic and structure priors and how we use the priors to the denoised U-Net in detail.
III-A Overview
In this paper, we propose a diffusion-based realistic image restoration method with semantic and structure priors. We fully leverage the generative priors of pre-trained T2I models such as stable diffusion to generate realistic results. However, due to the uncontrollability of the diffusion process, we utilize semantic and structure information from low-quality images as guidance, aiming to generate realistic restoration results with faithful structures and clear details. The overall framework of our model is depicted in Fig 2. We design specific modules to extract semantic and structure priors. The semantic and structure priors are integrated into the denoising U-Net through Prior-Guided Attention modules.
We design the Semantic Prior extraction module with two branches. The first branch utilizes the inference capability of a pre-trained multimodal language model to generate text embedding, which is explicit and can be modified by the user to suit their needs. The second branch utilizes the CLIP image encoder and a Refine Layer to generate the image embedding, which is implicit and contains the semantic information that cannot be expressed by text. The explicit text embedding and implicit image embedding collaborate as the semantic prior for the diffusion model’s control.
To get the degradation-independent Structure Prior, we employ a constrained Processor and a ControlNet module. The Processor extracts degradation-independent features from the low-quality image under the supervision of RGB loss and FFT loss at different scales, while the ControlNet module utilizes these features to generate structure control in the latent domain.
Our model is trained end-to-end, unlike SeeSR [34] and CoSeR [35], which require additional training to obtain the semantic information from LQ images. With the aid of the powerful image understanding and reasoning capabilities of MLLM and CLIP, our method can directly extract meaningful semantic information from LQ images without the need for additional training.
III-B Explicit and Implicit Semantic Prior.
III-B1 Analysis of the Influence of Semantic Prior
Semantic Prompts are key factors in determining the generation results of generative models. Therefore, how to mine semantic priors from low-quality images is an important issue we should consider. We use PASD [33] as a baseline and employ null prompts, LQ prompts, and GT prompts to guide image restoration individually. The results are shown in Fig 3. We have two observations:
  • (1)
    Incorporating textual prompts contributes to generating details in images. The more details the prompts provide, the richer details will be generated.
  • (2)
    Due to the influence of image degradation and the limitation of BLIP’s caption generation capability, directly using the LQ prompts for image generation can lead to semantic errors (e.g., as shown in Fig 3 (b), where an airplane is incorrectly generated).
Therefore, for realistic image restoration, we aim to infer reasonable and detailed semantic prompts from LQ images to guide the generation process. To get faithful and detailed semantic prompts, we propose a collaboration strategy to get explicit and implicit semantic prior from LQ images.